OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8=y A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3)

OpenStudy (anonymous):

\[ y=x^2-6x+8\\ y'= 2 x -6\\ 2 x -6 =0\\ x=3\\ y= 9 -18 +8=-1\\ (3,-1) \]

OpenStudy (anonymous):

Now do it by completing the square.

OpenStudy (anonymous):

\[ y=x^2-6x+8=y= x^2 - 6 x + 9 -9 + 8= (x-3)^2 -1 \]